Directrix (directrix):

Digits Problem: Would you look at the solution for the following problem and comment on whether it is correct? Thanks. Problem appears first in thread followed by proposed solution.

Directrix (directrix):

If the six digits 1, 2, 3, 5, 5, and 8 are randomly arranged into a six digit positive integer (number), what is the probability that the integer (number) is divisible by 15? Write your answer as a fraction in lowest terms.

Directrix (directrix):

To be divisible by 15 a number has to be divisible by 3 and by 5. To be divisible by 5 the number must end in a 0 or a 5. However, divisibility by 3 is not as simple - the rule is to add up all of the digits of the number and if the sum is a multiple of 3 then the original number is also --------------------------- The sum of these digits 1, 2, 3, 5, 5, and 8 is 24 which is a multiple of 3. Of the six digits, two are the number 5 and there are no 0s. For the last digit, there are two choices: 5 or 5. After using one of these two choices for the last digit, there are five choices left for the remaining 5 digits: 1,2,3,5,8. There are 5*4*3*2*1*2 = 240 ways to create a number divisible by 15 from the digits. The two fives are indistinguishable. Hence, there are 240/2 = 120 ways to write distinct numbers that adhere to the requirements. 120 will be the numerator of the probabiity fraction. For the six digits, there are: 6*5*4*3*2*1 = 720 possible numbers. Because of the two fives, there are 360 distinct possible numbers. P(the integer (number) is divisible by 15) = 120/360 = 1/3.

OpenStudy (kinggeorge):

Looks correct and makes sense to me.

Directrix (directrix):

If I don't divide out the repetitions for the numerator and denominator, the probability answer is the same. I thought about not mentioning it in the explanation for the student. But, somehow, I am thinking the duplications have to matter but not here, it seems. @PaxPolaris

OpenStudy (paxpolaris):

you're right it's still 1/3

Directrix (directrix):

Then, I was looking at the following problem: How many different messages can be represented by sequences of length 3 that use two dashes and one dot? and thinking 3*2*1 = 6 but because of the two indistinguishable dashes, then the answer should be 3. Does anybody agree?

OpenStudy (kinggeorge):

I agree. Possibilities are .-- -.- --.

OpenStudy (anonymous):

That is corect. Even the probability thing. Any repitions - You divide their interarrangeent sit doesnt matter. So if n similar things are thee in arranements you divide by n!.